Two incinerators are being considered by a waste management company. Design A has an initial cost of $2,500,000, has annual operating and maintenance costs of $800,000, and requires overhauls every 5 years at a cost of $1,250,000. Design B is more sophisticated, including computer controls; it has an initial cost of $5,750,000, has annual operating and maintenance costs of $600,000, and requires overhauls every 10 years at a cost of $3,000,000. Using a 5% per year interest year, design the capitalized cost for each design and recommend which should be chosen.
Answer #1

Design A=2500000+800000/5%+1250000/1.05^5*1/(1-1/1.05^5)=$23,024,369.95

Design B=5750000+600000/5%+3000000/1.05^10*(1-1/1.05^10)=$18,461,071.31

Choose Deisgn B as it has lower capitalized cost